# Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study over Open-ended Question Answering

### v1

@@ A.1 Prompt List
A.1.1 Knowledge-Augmented Language Model Prompting
System Instruction: “You are a helpful assistant designed to answer the users’ open-ended questions. Your task is to provide accurate, concise, and useful information to foster understanding and solve problems. Whether the questions relate to complex scientific concepts, historical events, practical advice, or everyday life, your goal is to assist by offering thoughtful and informative responses."
Example 1:
Question: How do you think blockchain technology can be applied to secure online voting systems? Please provide specific suggestions?
Below are the facts that might be relevant to answer the question: (’Bitcoin’, ’wikiPageWikiLink’, ’Bill_Gates’), (’Bitcoin’, ’wikiPageWikiLink’, ’Non-fungible_token’), (’Decentralized_application’, ’wikiPageWikiLink’, ’Cardano_(blockchain_platform)’), (’Non-fungible_token’, ’wikiPageWikiLink’, ’Decentralized_application’), (’Hyperledger’, ’wikiPageWikiLink’, ’Blockchain’), (’Blockchain’, ’implementations’, ’Bitcoin’), (’Blockchain’, ’currencyName’, ’Litecoin’), (’Thing’, ’type’, ’Privacy_and_blockchain’), (’Thing’, ’type’, ’Saraju_Mohanty’), (’Thing’, ’type’, ’Blockchain’), (’Thing’, ’type’, ’Satoshi_Nakamoto’), (’Thing’, ’type’, ’Pseudonym’), (’Thing’, ’type’, ’Metaverse’), (’Thing’, ’type’, ’PricewaterhouseCoopers’)
Answer: Blockchain technology can significantly enhance the security of online voting systems through its decentralized, transparent, and immutable nature. By distributing data across multiple nodes, blockchain eliminates single points of failure, reducing the risk of hacking or vote tampering. The system allows for verifiable, yet anonymous, vote recording through cryptographic methods, ensuring voter privacy while preventing fraud such as double voting. Additionally, once recorded, votes cannot be altered, safeguarding election integrity. Smart contracts can further automate the voting process, enabling instant and accurate tallying, while the system’s resilience to cyberattacks like DDoS ensures uninterrupted functionality.
Other In-Context Few-shots
@@ A.1.1 Knowledge-Augmented Language Model Prompting
A.1.2 OKGQA Query Generation Prompting
System Instruction: Generate open-ended questions about different types: character description, event description, cause explanation, relationship explanation, trend prediction, outcome prediction, contrast analysis, historical comparison, evaluation and reflection, and application and practice.
Some templates are provided below: character description: describe a [person]’s significant contributions during their career. Example: Please describe Albert Einstein’s contributions to the field of physics. event description: Provide a detailed description of the background and course of an [event]. Example: Please provide a detailed description of the background and course of the French Revolution. cause explanation: Why did [person] take [action] at [time]? Example: Why did Nixon choose to resign from the presidency in 1974? relationship explanation: Explain the relationship between [entity A] and [entity B] and its significance. Example: Explain the relationship between Alexander the Great and Aristotle and its significance. trend prediction: Based on the historical behavior of [entity], what do you think it might do in the future? Example: Based on Tesla’s historical behavior, in which fields do you think it might innovate in the future? outcome prediction: Based on the current situation, how do you predict [event] will develop? Example: Based on the current international situation, how do you predict climate change policies will develop? contrast analysis: Compare and contrast the similarities and differences between [entity A] and [entity B] in [aspect]. Example: Compare and contrast the leadership styles of Steve Jobs and Bill Gates. historical comparison: Compare the impact of [historical event A] and [historical event B]. Example: Compare the impact of World War I and World War II on the global order evaluation and reflection: How do you evaluate the impact of [person/event] on [field]? Please explain your viewpoint. Example: How do you evaluate Martin Luther King’s impact on the civil rights movement? Please explain your viewpoint. application and practice: How do you think [theory/method] can be applied to [practical issue]? Please provide specific suggestions. Example: How do you think machine learning technology can be applied to medical diagnostics? Please provide specific suggestions. Generate the questions, the type of the questions, the placeholders, the naturalness of your generated questions (choose from high, medium, and unnatural), the difficulty of the generated questions (choose from hard, medium and easy) and dbpedia_entities (link the placeholders to dbpedia entities) in JSON format.
@@ A.1.2 OKGQA Query Generation Prompting
A.2 Prompts for Instruction Tuner
Act as an "Instruction Tuner" for the LLM, you will be given the inputs: (1) the {Current Instruction} used to guide the LLMs’s evaluation, including specific examples with ground truth labels; (2) {Current Errors} that emerged with this instruction was applied to the dataset.
@@ A.2 Prompts for Instruction Tuner
A.3 Metrics Prompt for G-eval
@@
System Instruction: “You are a helpful assistant designed to evaluate the quality of the response to a query. Your task is to rate the response on one metric defined as below:"
@@
Empowerment Criteria: Evaluate whether the ‘Actual Output’ can help the reader understand the topic and make informed decisions regarding the ‘Input’. A response with high empowerment provides accurate information and explanations that enhance the reader’s understanding. When evaluating empowerment, consider the relevance of the information provided in the ‘Actual Output’ to the ’Input’ and the ’Retrieval Context’.
@@
Comprehensiveness Criteria: Evalute the extent to which the ’Actual Output’ covers all aspects and details of the question ’Input’. A comprehensive answer should thoroughly address every part of the question, leaving no important points unaddressed. When evaluating comprehensiveness, consider the relevance of the information provided in the ’Actual Output’ to the ’Input’ and the ’Retrieval Context’.
@@
Correctness Criteria: Measure how clearly and specifically the ’Actual output’ responds to the question ’input’. A highly direct response stays focused on the question, providing clear and unambiguous information. When evaluating correctness, consider the relevance of the information provided in the ’Actual Output’ to the ’Input’ and the ’Retrieval Context’.
@@
Context Relevance Criteria: Evaluate the extent to which the ’Actual output’ incorporates relevant information from the ’Retrieval Context’. This includes assessing whether the output adheres to the thematic, factual, and situational specifics presented in the ’Retrieval Context’. Relevant responses not only address the direct query but also align closely with the contextual elements provided, ensuring a seamless and coherent transition between the ’Retrieval Context’ and the ’Actual Output’. The most contextually relevant responses demonstrate an understanding and appropriate reflection of the given circumstances, historical facts, or conceptual background, thereby contributing to the overall accuracy and utility of the information provided.
A.4 Query Construction
In this section, we discuss the details of the query construction of OKGQA. It follows a template-based method with LLMs to generate a diverse range of queries. To ensure that the generated queries accurately represent real-world scenarios and complexities, we integrate the corresponding Wikipedia pages of each entity as contextual information for the LLMs’ generation. Furthermore, we follow a human-in-the-loop process to optimize the instruction used for generation, as shown in Figure 6. Intuitively, starting with an initial instruction, we generate a large number of query candidates, then use an LLM as automatic evaluator to evaluate the quality of the query, denoted as a set of scores from different metrics $s_{\text{auto}}$ ranging from 1 to 10, with higher scores indicating better performance. Then, these queries are manually evaluated and assign a set of human-label score $s_{\text{human}}$ correspond to $s_{\text{auto}}$, normalized as the same range as $s_{\text{auto}}$.We then optimize the input instructions by iteratively generate instructions to minimize the gap between $s_{\text{human}}$ and $s_{\text{auto}}$. This process is quite mimic the way of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and inherits the benefit that labeling the reward or penalty of the LLMs’ output is much easier than labeling the output directly.
Figure 6: Process of human-in-the-loop of query construction.
A.5 Personalized PageRank (PPR)
In this section, we discuss the details of the PPR algorithm used in Section 1 to prune the graph from DBPedia and concentrate on nodes most pertinent to the central nodes of interest. The PPR is calculated using the iterative formula:
$\mathbf{p}=\alpha\mathbf{A}^{\top}\mathbf{p}+(1-\alpha)\mathbf{s},$ (3)
[…]
A.6 Prize-Cost-based Path Retrieval
The path-retrieval method is designed to construct and evaluate paths in a graph based on predefined prize assignments and cost allocations. The objective is to form sequences of nodes and edges, represented as $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$, that maximize the overall score and minimize the costs. To efficiently manage the exploration of potential paths, we utilizes a priority queue, a data structure that allows paths to be organized based on their scores, ensuring that the highest-scoring paths are processed first. The method starts by picking a number of starting nodes with high prizes. We then expand these starting points by exploring neighboring nodes. For each neighbor, the method calculates a new score. This score is the sum of the neighbor’s prize and the edge’s prize minus the edge’s cost. If this neighbor hasn’t been visited before, which helps avoid looping, the algorithm adds this neighbor to the path. This new path is then added to the priority queue. This expansion keeps going until the path reaches a maximum length or can’t be extended further. The algorithm keeps track of paths already explored to avoid repetition and ensure paths don’t loop back on themselves. When no more paths can be added or the priority queue is empty, the algorithm sorts the paths by their scores from highest to lowest.
A.7 KG Similarity Metrics
To measure how much the perturbed KGs are deviated from the original KG used in Section 3. We re-present the according metrics from (Raman et al., 2020) below:
Similarity in Degree Distribution (SD2)  SD2 also measures structural similarity between two KGs, while addressing SC2D’s ineffectiveness when the KGs’ entities have tiny local clustering coefficients (e.g., the item KG used by recommender systems is roughly bipartite). In such cases, SC2D is always close to one regardless of perturbation method, thus rendering SC2D useless. Let $\mathbf{d}^{r}$ denote the $|\mathcal{E}|$-dimensional degree vector for $\mathcal{G}^{r}$, where the $i$th element of $\mathbf{d}^{r}$ is $\text{deg}(e_{i})$. Then, the mean degree vectors for $\mathcal{G}$ and $\mathcal{G}^{\prime}$ are $\mathbf{d}_{o}=\frac{1}{|\mathcal{R}|}\sum_{r\in\mathcal{R}}\mathbf{d}^{r}$ and $\mathbf{d}_{p}=\frac{1}{|\mathcal{R}^{\prime}|}\sum_{r\in\mathcal{R}^{\prime}}\mathbf{d}^{r}$, respectively. SD2 is defined as $f_{\text{SD2}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=1-\frac{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}}{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}+1}\in[0,1]$, with higher value indicating higher similarity.

### v2

@@ Appendix C Prompt List
C.1 Knowledge-Augmented Generation
System Instruction: “You are a helpful assistant designed to answer the users’ open-ended questions. Your task is to provide accurate, concise, and useful information to foster understanding and solve problems. Whether the questions relate to complex scientific concepts, historical events, practical advice, or everyday life, your goal is to assist by offering thoughtful and informative responses.”
In-Context Few-shots
@@ C.1 Knowledge-Augmented Generation
C.2 OKGQA Query Generation Prompting
System Instruction: “Generate open-ended questions about different types: character description, event description, cause explanation, relationship explanation, trend prediction, outcome prediction, contrast analysis, historical comparison, evaluation and reflection, and application and practice. Please provide specific suggestions. Generate the questions, the type of the questions, the placeholders, the naturalness of your generated questions (choose from high, medium, and unnatural), the difficulty of the generated questions (choose from hard, medium and easy) and DBPedia_entities (link the placeholders to DBPedia entities) in JSON format.”
Example 1: as shown in Figure 8.
Figure 8: Example 1 Demonstration.
@@ C.2 OKGQA Query Generation Prompting
C.3 Prompts for Instruction Tuner
Act as an “Instruction Tuner” for the LLM, you will be given the inputs: (1) the {Current Instruction} used to guide the LLMs’s evaluation, including specific examples with ground truth labels; (2) {Current Errors} that emerged with this instruction are applied to the dataset.
@@ C.3 Prompts for Instruction Tuner
C.4 Metrics Prompt for G-eval
@@
System Instruction: “You are a helpful assistant designed to evaluate the quality of the response to a query. Your task is to rate the response on one metric defined as below:”
@@
Empowerment Criteria: Evaluate whether the “Actual Output” can help the reader understand the topic and make informed decisions regarding the “Input”. A response with high empowerment provides accurate information and explanations that enhance the reader’s understanding. When evaluating empowerment, consider the relevance of the information provided in the “Actual Output” to the “Input” and the “Retrieval Context”.
@@
Comprehensiveness Criteria: Evaluate the extent to which the “Actual Output” covers all aspects and details of the question “Input”. A comprehensive answer should thoroughly address every part of the question, leaving no important points unaddressed. When evaluating comprehensiveness, consider the relevance of the information provided in the “Actual Output” to the “Input” and the “Retrieval Context”.
@@
Correctness Criteria: Measure how clearly and specifically the “Actual output” responds to the question “input”. A highly direct response stays focused on the question, providing clear and unambiguous information. When evaluating correctness, consider the relevance of the information provided in the “Actual Output” to the “Input” and the “Retrieval Context”.
@@
Context Relevance Criteria: Evaluate the extent to which the “Actual output” incorporates relevant information from the “Retrieval Context”. This includes assessing whether the output adheres to the thematic, factual, and situational specifics presented in the “Retrieval Context”. Relevant responses not only address the direct query but also align closely with the contextual elements provided, ensuring a seamless and coherent transition between the “Retrieval Context” and the “Actual Output”. The most contextually relevant responses demonstrate an understanding and appropriate reflection of the given circumstances, historical facts, or conceptual background, thereby contributing to the overall accuracy and utility of the information provided.
Response: [Respond with metric and the corresponding score.]